With what velocity must an object. Be thrown vertically upward or authority in the maximum height of 900ft from its starting point.

1
Expert's answer
2021-11-26T14:46:03-0500

"v^2=u^2-2gh"

But u=0

"v^2=-2gh"

h=900ft=274.32m

"\therefore v^2=-2*-9.81*274.32"

"v^2=5382.1584"

"v=\sqrt{5382.1584}"

"v=73.36m\/s"